A different derby car has a mass of 170 kg. If the car traveled across a 10 meter rough road before it came to a stop. How much work did the road do to stop the car? (Assume the car started with 13m/s)
1
Expert's answer
2020-06-02T17:52:49-0400

"W=KE=\frac{mv^2}{2}=\frac{170\cdot 13^2}{2}=14365J"